OpenStudy (anonymous):

A heavy load carrier travels 360 miles through the flatlands in the same amount of time that it travels 200miles in the mountains. The speed of truck is 20miles per hour faster on level ground hen it's speed through the mountains. find the speed of the truck in both the flatlands and the mountains?

OpenStudy (anonymous):

s - speed on the moutain s + 20 - speed on flatland distance = speed x time 360 miles = (s+20)* t 200 miles = s *t divide first equation to second equation 360/ 200 = (s+20)/s 360S = (S+20) 200 360S = 200 S + 4000 160S = 4000 S = 4000/ 160 <=== divide to get speed on moutain Then take the result of S plus 20 to get speed on flatland

OpenStudy (anonymous):

so when i divide 4000 by 160 i got 87.5 and then i would take 87.5+20 and get the speed on flatland?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

to test it...take 360/speed of flatland and take 200/ speed of moutain...if they came out the same time result...we did it right

OpenStudy (dumbcow):

4000/160 = 25

OpenStudy (anonymous):

ooh ya my bad thnks